### v1

@@ 5.4 Variations
We could use a different definition of “small” for our sets $S_{\alpha}(\mu)$, now in the sense of the overcoverage probability. However, this approach can lead to very asymmetric results without some additional constraints. We could combine the idea with the minimal number of observations of section 5.2. We look for the smallest number of contiguous observations satisfying the probabilty content, selecting the end points that minimize the overcoverage. That is, we relax the strict ordering on probability of Sterne but also do not require the minimal upper bound on the confidence interval property of Crow&Gardner. There are many other variations of these general ideas we could try. Indeed, it is possible to improve coverage. However, the intervals then tend to be large, and less symmetric in interval, to the point where such results do not seem attractive.
@@ 5.5 Kabaila-Byrne intervals
Kabaila and Byrne [23] propose a confidence interval that is “shortest” in the sense that any further shortening would cause the interval to undercover for some values of $\mu$. More completely, their interval is designed to satisfy the criteria of optimal coverage, shortest length, and strict ordering. The reader is referred to [23] for the algorithm, which we have implemented in MATLAB. Fig. 9 shows a comparison of Kabaila&Byrne and Crow&Gardner confidence intervals. They generally track, but the Kabaila&Byrne intervals tend to be shifted higher and are longer intervals than Crow&Gardner.
@@ 5.6 Inverting a likelihood ratio test
Now we build the acceptance region for the test. If the likelihood ratio is large, the evidence favors $H_{0}$. Thus, we will include in the acceptance region those values of $N=n$ for which the likelihood ratio is large. We’ll stop adding values of $n$ when we have reached the desired probability. That is, we have acceptance region
@@
We can perform this procedure for any desired value of $\mu=\mu_{0}$. We could check whether this procedure results in sets of contiguous values of $n$. However, we’ll adopt an algorithm that guarantees it, at the cost of possibly not strictly ordering on the ratio for some values of $\mu$.
@@
Because this method is based on the likelihood ratio, we denote it as the “likelihood ratio” (LR) method. Fig. 10 compares the properties of the LR method with Crow&Gardner.
@@ 5.7 Inverting a score test
We define our test statistic as follows: The score function for a parameter $\theta$ and observation $x$ is
@@ 6 Square root of NN
The coverage probability of the $\sqrt{n}$ intervals is shown as a function of $\mu$ in Fig. 12. As expected, the intervals both undercover and overcover, depending on the value of $\mu$, hence don’t meet our coverage requirement. However, at least for large enough $\mu$ the approximation might be good enough for certain purposes, such as graphical representation.
@@ 7 Frequency properties of Bayes’ intervals with uninformative priors
As mentioned in the introduction, we may examine the frequency properties of objective Bayes methods. For illustration and comparison, we consider two popular choices of “uninformative” reference prior, the uniform prior and the Jeffreys prior. Here, we need to explicitly include the background parameter.
@@
where $P(\theta;b)$ is the prior distribution. The Bayes interval for a given confidence level $1-\alpha$ is then determined by including all values of $\theta\in(\theta_{\rm min},\theta_{\rm max})$ such that
@@ 7.2 Jeffreys prior
For our Poisson distribution with background $b$, this works out to [^10] One could also consider a prior proportional to $1/\sqrt{\theta}$, but this is not the Jeffreys prior unless $b=0$.
@@
again assuming on physical grounds that $\theta\geq 0$. The posterior distribution is thus
@@
Figure 13 shows a comparison of the Bayes 68% confidence credible intervals for the uniform and Jeffreys priors, for a background parameter $b=5$. They are similar, but the preference for $\theta=0$ of the Jeffreys prior is seen in the lower values of the interval boundaries and the smaller interval sizes. The coverage is observed to both under and over cover, depending on $\theta$. This is unsurprising, as there is no requirement on coverage for Bayes intervals. The Bayes intervals also tend to be shorter in length than $2\sqrt{n}$ for small $n$, unlike the behavior of the frequency intervals.
Figure 13: Bayes 68% confidence credible intervals for $b=5$, comparing uniform and Jeffreys priors. Left: coverage probability as a function of $\theta$; Middle: confidence interval boundaries as a function of observed counts. Circles are for uniform prior, squares for Jeffreys prior. The plus symbols show the maximum likelihood estimator for $\theta$; Right: size of the confidence interval as a function of observed counts. Circles are for uniform prior, squares for Jeffreys prior. The plus symbols show the values of $2\sqrt{n}$.
@@
Figure 14: Comparison of Crow&Gardner 68% confidence intervals and Bayes 68% confidence credible intervals (with a uniform prior) for $b=5$. Left: coverage probability as a function of $\theta$; Middle: confidence interval boundaries as a function of observed counts. Circles are for Crow&Gardner, squares for uniform prior. The plus symbols show the maximum likelihood estimator for $\theta$; Right: size of the confidence interval as a function of observed counts. Circles are for Crow&Gardner, squares for uniform prior. The plus symbols show the values of $2\sqrt{n}$.
@@ 8.1 The $CL_{s}$ method
Note that this isn’t the same test as we considered in Eq. 47, since $\theta$ is specified under the alternative. The likelihood ratio is uniformly most powerful for a simple test, so it is natural to use here. We thus take as our test statistic, the likelihood ratio between the two hypotheses, [^11] We define this here to be consistent with the convention for the $CL_{s}$ method and attach a prime to flag the difference in convention with Eq. 48.
@@
This ratio is denoted by $Q$ in papers on the $CL_{s}$ method. Here we have possible values $0<\lambda^{\prime}<\infty$. We generally work with $R\equiv-2\log\lambda^{\prime}$, with range $(-\infty,\infty)$, where negative values favor $H_{1}$ and positive values favor $H_{0}$.
@@
$\begin{split}1-p_{0}(n)&\equiv P\left[R(N)\geq R(n)|H_{0}\right]\\ 1-p_{1}(n)&\equiv P\left[R(N)\geq R(n)|H_{1}\right].\end{split}$ (77)
We may note that the quantity $1-p_{1}$ is simply the $p$ value for hypothesis $H_{1}$. The quantity
$p_{0}(n)=P\left[R(N)<R(n)|H_{0}\right]$ (78)
is the $p$ value for $H_{0}$ in our hypothesis test, except for the non-zero measure at $N=n$ according to the discreteness of the distribution.
@@
The idea of dividing by $1-p_{0}$ protects against excluding $\theta_{1}$ in the following situation: Suppose $\theta_{0}$ and $\theta_{1}$ are not so different, that is the possible signal is “close” to zero. Suppose the $p$ value for $\theta_{1}$ is very small, that is, it provides a poor fit to the data. In this case, $H_{0}$ and $H_{1}$ are about the same, so $1-p_{0}$ and $1-p_{1}$ will be nearly the same. Hence $CL_{s}\approx 1$, and we do not exclude the possibility of a small signal. That is, we are protected against the possibility of excluding the presence of a small signal even if the fit is poor. We see that the $CL_{s}$ method is designed to provide “limits”, never excluding the possibility that there is no signal. This method, motivated by the desire to avoid incorrect statements about the parameter, clearly goes beyond simply description. If a fit to the data is poor, that is an important descriptive statement.
We see that $CL_{s}$ is generally greater than $1-p_{1}$, the $p$ value for the signal hypothesis. Thus, an interval based on $CL_{s}(n;\theta_{1})<\alpha$ is conservative – the exclusion region corresponds to probability smaller than $\alpha$.
We now apply the $CL_{s}$ method to our Poisson problem. We have $\theta_{0}=0$. Then
@@
$\begin{split}1-p_{0}&=P(N\leq n|H_{0})\\ 1-p_{1}&=P(N\leq n|H_{1}).\end{split}$ (81)
@@
The $CL_{s}$ limit converges on the standard upper limit of section 4 in the limit of no background. For $b>0$, however, the two limits are not the same. The largest difference, for a given background is that the $CL_{s}$ limit tends to overcover substantially more than the standard limit for small values of $\theta$, as seen for $b=5$ in Fig. 15. This is consistent with our observation above that the division by $1-p_{0}$ leads to a conservative upper limit.
Figure 15: Performance of 90% confidence upper limits, comparing $CL_{s}$ with the standard upper limit for $b=5$. Left: coverage probability as a function of $\theta$; Right: upper limit as a function of observed counts. Circles are for $CL_{s}$, squares for standard upper limit.
@@ 8.2 Feldman-Cousins
As mentioned above, if there is no background the FC and LR intervals are the same, because Eqn. 48 and Eqn. 83 are identical for zero background. For a background of $b=5$, the comparison between the properties of the FC and LR intervals is shown in Fig. 16. With background, there are substantial differences between the two intervals. The coverage is very similar between the two, except for small values of $\theta$, where the FC coverage is greater than the LR, due to the interval being forced positive for small $n$. The intervals are very different for small observations. The FC intervals, by design, do not include negative values, while the LR intervals can include negative values, reflecting occurrences of negative fluctuations in the background. For small values of observation $n$, the FC intervals are considerably shorter than the LR intervals.
@@
Figure 16: Performance of 68% confidence intervals for $\theta$, comparing FC with the likelihood ratio intervals for $b=5$. Left: coverage probability as a function of $\theta$; Middle: confidence interval as a function of observed counts. Circles are for LR, squares for FC. The plus symbols show the maximum likelihood estimator for $\theta$; Right: interval size as a function of observed counts. Circles are for LR, squares for FC. The plus symbols show the values of $2\sqrt{n}$.
@@
However, we may observe that the two intervals are not so different in the following sense: The coverage (and hence the validity of the interval as a frequency confidence interval) is not affected if we simply cut the LR interval at $\theta=0$, since we are assuming that the true value of $\theta$ must be non-negative. We’ll refer to these intervals as cut LR intervals. Note that this procedure can yield intervals of zero size, when a sufficiently large negative fluctuation has occurred. For the graphical indication of this, we’ll plot this as upper and lower bounds equal to zero, with the understanding that it is a null interval. The comparison with FC, again for $b=5$, is shown in Fig. 17. As expected, the coverage is identical with Fig. 16. However, the intervals and their sizes are now very similar between the FC and cut LR methods, on the visual scale of the graphs. When the cut LR intervals are very small, or null (occurring when a large negative fluctuation has occurred), the FC interval is also very small, though by design never null.
@@
Figure 17: Performance of 68% confidence intervals for $\theta$, comparing FC with the likelihood ratio intervals for $b=5$. The likelihood ratio intervals have been cut so that only the positive portion of the interval is retained. Left: coverage probability as a function of $\theta$; Middle: confidence interval as a function of observed counts. Circles are for LR, squares for FC. The plus symbols show the maximum likelihood estimator for $\theta$; Right: interval size as a function of observed counts. Circles are for LR, squares for FC. The plus symbols show the values of $2\sqrt{n}$.
@@
What do we make of this? The (uncut) LR interval has similar behavior to the other conventional methods of Section 5, differing in relatively small ways in performance. In particular, all of these intervals will become negative in the presence of large downward background fluctuations. As we have discussed, this is quite natural and provides an easy-to-understand description of the measurement. When we cut it off, we don’t lose frequency validity, but the value as a description of the measurement is diminished. Null intervals may be obtained for multiple possible outcomes, so just how much background (plus signal) fluctuation has occurred has been lost. Even when the interval is not completely null, the precision of the measurement is obscured.
@@
The FC interval was designed, in large part at least, to overcome the tendency of physicists to change their analysis depending on the outcome of the measurement, which leads to biased results (in the frequency sense). In particular, the issue addressed is the tendency to decide whether to quote a central interval or a one-sided interval depending on the observation. The FC interval “makes this decision” for you and preserves the strict frequency interpretation. Of course, this isn’t the only way to accomplish this. [^13] But far better is to design the analysis ahead of time so that such bias is avoided. Any of the conventional methods of Section 5 will also do this, if we simply exclude any negative values from the interval. It could be remarked that there is a distinction, because the FC intervals are never empty, while simply cutting out negative values in the other methods leads to null intervals for some observations. However, these can also readily be made to be non-empty if that is desired. For example, one way to do this is, whenever an empty interval is encountered, to simply modify it to an interval $0,\epsilon]$, where $\epsilon$ is a small positive number. This could be done in a way that maintains sufficiency, e.g., $\epsilon=1/[\hbox{large number}\times(1+n)]$. This is of course contrived, but that is really the point, attempting to restrict description to non-negative intervals leads to cryptic results.

### v2

@@ 6.4 Variations
We could use a different definition of “small” for our sets $S_{\alpha}(\mu)$, now in the sense of the overcoverage probability. However, this approach can lead to very asymmetric results without some additional constraints. We could combine the idea with the minimal number of observations of section 6.2. We look for the smallest number of contiguous observations satisfying the probability content, selecting the end points that minimize the overcoverage. That is, we relax the strict ordering on probability of Sterne but also do not require the minimal upper bound on the confidence interval property of Crow&Gardner. There are many other variations of these general ideas we could try. Indeed, it is possible to improve coverage, for example by working farther into the tails of the distribution. However, the intervals then tend to be large, and less symmetric in interval, to the point where such results do not seem attractive.
@@ 6.5 Kabaila-Byrne intervals
Kabaila and Byrne [20] propose a confidence interval that is “shortest” in the sense that any further shortening would cause the interval to undercover for some values of $\mu$. More completely, their interval is designed to satisfy the criteria of optimal coverage, shortest length, and strict ordering. The reader is referred to [20] for the algorithm, which we have implemented in MATLAB. Fig. 9 shows a comparison of Kabaila&Byrne and Crow&Gardner confidence intervals. They generally track, but the Kabaila&Byrne intervals tend to be shifted higher and are longer intervals than Crow&Gardner. However, the Kabaila&Byrne intervals are constructed requiring strict ordering, which is not satisfied by Crow&Gardner intervals [20, 37].
@@ 6.6 Inverting a likelihood ratio test
Now we build the acceptance region for the test. If the likelihood ratio is large, the evidence favors $H_{0}$. Thus, we will include in the acceptance region those values of $N=n$ for which the likelihood ratio is large. We will stop adding values of $n$ when we have reached the desired probability. That is, we have acceptance region
@@
We can perform this procedure for any desired value of $\mu=\mu_{0}$. We could check whether this procedure results in sets of contiguous values of $n$. However, we will adopt an algorithm that guarantees it, at the cost of possibly not strictly ordering on the ratio for some values of $\mu$.
@@
Because this method is based on the likelihood ratio, we denote it as the “likelihood ratio” (LR) method. Fig. 10 compares the properties of the LR method with Crow&Gardner. We see that the coverage for LR intervals is often somewhat larger than for Crow&Gardner, and the intervals are also somewhat longer.
@@ 6.7 Inverting a score test
We define our test statistic as follows: In general, the score function for a parameter $\theta$ and observation $x$ is
@@ 7 Square root of NN
The coverage probability of the $\sqrt{n}$ intervals is shown as a function of $\mu$ in Fig. 12. As expected, the intervals both undercover and overcover, depending on the value of $\mu$, hence do not meet our coverage requirement. However, at least for large enough $\mu$ the approximation might be good enough for certain purposes, such as graphical representation.
@@ 8 Frequency properties of Bayes’ intervals with uninformative priors
As mentioned in the introduction, we may examine the frequency properties of objective Bayes methods. For illustration and comparison, we consider two popular choices commonly referred to as “uninformative” reference priors, the uniform prior and the Jeffreys prior. Here, we need to explicitly include the background parameter. We’ll assume that $\theta\geq 0$.
@@
where $P(\theta;b)$ is the prior distribution. The Bayes interval for a given probability level $1-\alpha$ is then determined by including all values of $\theta\in(\theta_{\rm min},\theta_{\rm max})$ such that
@@ 8.2 Jeffreys prior
For our Poisson distribution with background $b$, this works out to:
@@
again assuming on physical grounds that $\theta\geq 0$. We note that the prior depends on the background parameter, suggesting that it is not really capturing a degree-of-belief in $\theta$. [^12] One could also consider a prior proportional to $1/\sqrt{\theta}$, but this is not the Jeffreys prior unless $b=0$. In any event, we are looking here at descriptive properties of the method.
The posterior distribution is thus
@@
Figure 13 shows a comparison of the Bayes 68% confidence credible intervals for the uniform and Jeffreys priors, for a background parameter $b=5$. They are similar, but the preference for $\theta=0$ of the Jeffreys prior is seen in the lower values of the interval boundaries and the smaller interval sizes. The coverage is observed to both under and over cover, depending on $\theta$. This is unsurprising, as there is no requirement on coverage for Bayes intervals. The Bayes intervals also tend to be shorter in length than $2\sqrt{n}$ for small $n$, unlike the behavior of the conventional frequency intervals of section 6.
Figure 13: Bayes 68% probability credible intervals for $b=5$, comparing uniform and Jeffreys priors. Left: coverage probability as a function of $\theta$; Middle: confidence interval boundaries as a function of observed counts. Circles are for uniform prior, squares for Jeffreys prior. The plus symbols show the maximum likelihood estimator for $\theta$; Right: size of the confidence interval as a function of observed counts. Circles are for uniform prior, squares for Jeffreys prior. The plus symbols show the values of $2\sqrt{n}$.
@@
Figure 14: Comparison of Crow&Gardner 68% confidence intervals and Bayes 68% probability credible intervals (with a uniform prior) for $b=5$. Left: coverage probability as a function of $\theta$; Middle: confidence interval boundaries as a function of observed counts. Circles are for Crow&Gardner, squares for uniform prior. The plus symbols show the maximum likelihood estimator for $\theta$; Right: size of the confidence interval as a function of observed counts. Circles are for Crow&Gardner, squares for uniform prior. The plus symbols show the values of $2\sqrt{n}$.
@@ 9.1 The $CL_{s}$ method
Note that this is not the same test as we considered in Eq. 45, since $\theta$ is specified under the alternative. The likelihood ratio is uniformly most powerful for a simple test, so it is natural to use here. We thus take as our test statistic, the likelihood ratio between the two hypotheses, [^13] We define this here to be consistent with the convention for the $CL_{s}$ method and attach a prime to flag the difference in convention with Eq. 46.
@@
This ratio is denoted as $Q$ in papers on the $CL_{s}$ method. Here we have possible values $0<\lambda^{\prime}<\infty$. We generally work with $R\equiv-2\log\lambda^{\prime}$, with range $(-\infty,\infty)$, where negative values favor $H_{1}$ and positive values favor $H_{0}$.
@@
$\begin{split}1-p_{0}(n)&\equiv\hbox{Prob}\left[R(N)\geq R(n)|H_{0}\right]\\ 1-p_{1}(n)&\equiv\hbox{Prob}\left[R(N)\geq R(n)|H_{1}\right].\end{split}$ (75)
We may note that the quantity $1-p_{1}$ is simply the $p$-value for hypothesis $H_{1}$. The quantity
$p_{0}(n)=\hbox{Prob}\left[R(N)<R(n)|H_{0}\right]$ (76)
is the $p$-value for $H_{0}$ in our hypothesis test, except for the non-zero measure at $N=n$ according to the discreteness of the distribution.
@@
The idea of dividing by $1-p_{0}$ protects against excluding $\theta_{1}$ in the following situation: Suppose $\theta_{0}$ and $\theta_{1}$ are not so different, that is the possible signal is “close” to zero. Suppose the $p$-value for $\theta_{1}$ is very small, that is, it provides a poor fit to the data. In this case, $H_{0}$ and $H_{1}$ are about the same, so $1-p_{0}$ and $1-p_{1}$ will be nearly the same. Hence $CL_{s}\approx 1$, and we do not exclude the possibility of a small signal. That is, we are protected against the possibility of excluding the presence of a small signal even if the fit is poor. We see that the $CL_{s}$ method is designed to provide “limits”, never excluding the possibility that there is no signal. This method, motivated by the desire to avoid incorrect statements about the parameter, clearly goes beyond simply description. If a fit to the data is poor, that is an important descriptive statement.
We see that $CL_{s}$ is generally greater than $1-p_{1}$, the $p$-value for the signal hypothesis. Thus, an interval based on $CL_{s}(n;\theta_{1})<\alpha$ is conservative – the exclusion region corresponds to probability smaller than $\alpha$.
We now apply the $CL_{s}$ method to our Poisson problem. We have $\theta_{0}=0$ in this case. Then
@@
$\begin{split}1-p_{0}&=\hbox{Prob}(N\leq n|H_{0})\\ 1-p_{1}&=\hbox{Prob}(N\leq n|H_{1}).\end{split}$ (79)
@@
The $CL_{s}$ limit converges on the standard upper limit of section 5 in the limit of no background. For $b>0$, however, the two limits are not the same. The largest difference, for a given background is that the $CL_{s}$ limit tends to overcover substantially more than the standard limit for small values of $\theta$, as seen for $b=5$ in Fig. 15. This is consistent with our observation above that the division by $1-p_{0}$ leads to a conservative upper limit. As a descriptive statement, the limit of section 5 is preferred.
Figure 15: Performance of 90% confidence upper limits, comparing $CL_{s}$ with the standard upper limit (section 5) for $b=5$. Left: coverage probability as a function of $\theta$; Right: upper limit as a function of observed counts. Circles are for $CL_{s}$, squares for standard upper limit.
@@ 9.2 Feldman-Cousins
As mentioned above, if there is no background the FC and LR intervals are the same, because Eqn. 46 and Eqn. 81 are identical for zero background. For a background of $b=5$, the comparison between the properties of the FC and LR intervals is shown in Fig. 16. With background, there are substantial differences between the two intervals. The coverage is very similar between the two, except for small values of $\theta$, where the FC coverage is greater than the LR, due to the interval being forced positive for small $n$. The intervals are very different for small observations. The FC intervals, by design, do not include negative values, while the LR intervals can include negative values, reflecting occurrences of negative fluctuations in the background. For small values of observation $n$, the FC intervals are considerably shorter than the LR intervals.
@@
Figure 16: Performance of 68% confidence intervals for $\theta$, comparing FC with the likelihood ratio intervals for $b=5$. Left: coverage probability as a function of $\theta$; Middle: confidence interval as a function of observed counts. Circles are for LR, squares for FC. The plus symbols show the maximum likelihood estimator for $\theta$ (allowed to be negative); Right: interval size as a function of observed counts. Circles are for LR, squares for FC. The plus symbols show the values of $2\sqrt{n}$.
@@
However, we may observe that the two intervals are not so different in the following sense: The coverage (and hence the validity of the interval as a frequency confidence interval) is not affected if we simply cut the LR interval at $\theta=0$, since we are assuming that the true value of $\theta$ must be non-negative. We will refer to these intervals as cut LR intervals. Note that this procedure can yield intervals of zero size, when a sufficiently large negative fluctuation has occurred. For the graphical indication of this, we will plot this as upper and lower bounds equal to zero, with the understanding that it is a null interval. The comparison with FC, again for $b=5$, is shown in Fig. 17. As expected, the coverage is identical with Fig. 16. However, the intervals and their sizes are now very similar between the FC and cut LR methods, on the visual scale of the graphs. When the cut LR intervals are very small, or null (occurring when a large negative fluctuation has occurred), the FC interval is also very small, though by design never null.
@@
Figure 17: Performance of 68% confidence intervals for $\theta$, comparing FC with the cut likelihood ratio intervals for $b=5$. The likelihood ratio intervals have been cut so that only the positive portion of the interval is retained. Left: coverage probability as a function of $\theta$; Middle: confidence interval as a function of observed counts. Circles are for cut LR, squares for FC. The plus symbols show the maximum likelihood estimator for $\theta$ (allowed to be negative); Right: interval size as a function of observed counts. Circles are for cut LR, squares for FC. The plus symbols show the values of $2\sqrt{n}$.
@@
What do we make of this? The (uncut) LR interval has similar behavior to the other conventional methods of Section 6, differing in relatively small ways in performance. In particular, all of these intervals will become negative in the presence of large downward background fluctuations. As we have discussed, this is quite natural and provides an easy-to-understand description of the measurement. When we cut it off, we do not lose frequency validity, but the value as a description of the measurement is diminished: Null intervals may be obtained for multiple possible outcomes, and how much background (plus signal) fluctuation has occurred has been lost. Even when the interval is not completely null, the precision of the measurement is obscured.
@@
The FC interval was designed, in large part at least, to overcome the tendency of physicists to change their analysis depending on the outcome of the measurement, which leads to biased results (in the frequency sense). In particular, the issue addressed is the tendency to decide whether to quote a central interval or a one-sided interval depending on the observation. The FC interval “makes this decision” for you and preserves the strict frequency interpretation. Of course, this is not the only way to accomplish this. [^15] But far better is to design the analysis ahead of time so that such bias is avoided. Any of the conventional methods of Section 6 will also do this, if we simply exclude any negative values from the interval. It could be remarked that there is a distinction, because the FC intervals are never empty, while simply cutting out negative values in the other methods leads to null intervals for some observations. However, these can also readily be made to be non-empty if that is desired. For example, one way to do this is, whenever an empty interval is encountered, to simply modify it to an interval $0,\epsilon]$, where $\epsilon$ is a small positive number. This could be done in a way that maintains sufficiency, e.g., $\epsilon=1/[\hbox{large number}\times(1+n)]$. This is of course contrived, but that is really the point, attempting to restrict description to non-negative intervals leads to cryptic results.
To be clear, I am not advocating use of the cut LR intervals even if modified to ensure no null intervals. The small intervals when the observation is small suggest greater precision than justified. The point is that the FC interval also suffers from this issue. FC [13] suggest to quote the “sensitivity” of the experiment to alert the consumer when this occurs.
We also compare the FC intervals with Bayes intervals with uniform prior in Fig. 18, for background $b=5$. For small observations $n$, the FC intervals are much shorter than the Bayes intervals, becoming somewhat larger at higher $n$. The comparison with Bayes intervals is consistent with the above observation about apparent precision.
Figure 18: Performance of 68% intervals for $\theta$, comparing FC confidence intervals with Bayes (uniform prior) credible intervals for $b=5$. Left: coverage probability as a function of $\theta$; Middle: confidence interval as a function of observed counts. Circles are for FC, squares for Bayes. The plus symbols show the maximum likelihood estimator for $\theta$ (allowed to be negative); Right: interval size as a function of observed counts. Circles are for FC, squares for Bayes. The plus symbols show the values of $2\sqrt{n}$.